OpenStudy (anonymous):

Find the product of (x + 3)(x − 3). x2 − 6x + 9 x2 + 6x + 9 x2 + 9 x2 − 9

OpenStudy (anonymous):

(x + 3)(x − 3)=x*x-3x+3x-3*3=x^2-9

OpenStudy (anonymous):

thx.